OpenStudy (anonymous):

Also cot(theta) if csc(theta)= 7/2 , how can you solve it

OpenStudy (unklerhaukus):

cot=1/tan tan=sin/cos csc=1/sin

OpenStudy (anonymous):

I mean "find the value of cot (x) if csc (x) = 3/2" i dont get it

OpenStudy (campbell_st):

well csc(x) 3/2 or 1/sin(x) = 3/2 therefore after taking the reciprocals sin(x) = 2/3 now draw a triangle  now cot = 1/tan so find tan(x) from the triangle above then find the reciprocal to find cot.